OpenStudy (anonymous):

Find the volume of the solid generated when the region under the curve y = 1/(x^2 + 3x + 2) from x = 0 to x = 1 is revolved about the given axis. Make certain to include sketches and justifications for each solid. a. x-axis b. y-axis

OpenStudy (amistre64):

youll most likely have to complete the square on the bottom and see if this changes into a 1/(u^2+1) format

OpenStudy (amistre64):

or do alot of partial decmomps

OpenStudy (amistre64):

decompositions

OpenStudy (anonymous):

(x+1)(x+2)

OpenStudy (amistre64):

hmmm, if you wanna go that route, partial decomp then

OpenStudy (amistre64):

either way, the radius is squared to find the area if a circle .... might wanna go ahead and plug a ^2 on those

OpenStudy (anonymous):

doesn't matter, that's the same thing

OpenStudy (amistre64):

\[\frac{1}{(x-a)^2(x-b)^2}=\frac{A}{x-a}+\frac{Bx+C}{(x-a)^2}+\frac{D}{x-b}+\frac{Ex+F}{(x-b)^2}\]

OpenStudy (anonymous):

i can integrate but my problem is, i can't make an equation

OpenStudy (amistre64):

when x=a \[1=\frac{D}{a-b}+\frac{Ea+F}{(a-b)^2}\] etc... gives you a system of equations to play with in the end

OpenStudy (amistre64):

you can make an equation?

OpenStudy (anonymous):

 like this